OpenStudy (anonymous):

Medal to help me solve and explain how to complete this broken down TY u+2=4

OpenStudy (eric_d):

rearrange it..

OpenStudy (eric_d):

put all the values to the left

OpenStudy (eric_d):

you'll find the unknown, u

OpenStudy (anonymous):

2? U+2=4 U+2=4 -2 -2 U=2

OpenStudy (anonymous):

why do they both turn into -2

OpenStudy (anonymous):

I always get mixed up with the signs

OpenStudy (anonymous):

You Have to Get the Variable All By itself U+2=4 -2 -2 If you are adding you have to subtract to both sides if you are Subtracting you add to both sides since you are adding you have to subtract To get U all by itself so 2-2=0 they cancel out 4-2=2 Since 2-2=0 now you have U all by itself so U=2 which you can prove by solving (2)+2=4

OpenStudy (eric_d):